# MENTOR: Mixture-of-Experts Network with Task-Oriented Perturbation for Visual Reinforcement Learning

### v1

@@ 3.2 Optimization: Task-oriented Perturbation Mechanism
Figure 5: Validation of task-oriented perturbation on Hopper Hop. Our method consistently achieves higher episode rewards with a consistent lower dormant ratio. The episode reward obtained by $\phi$ sampled from our method continuously increases and sometimes even surpasses that of the corresponding RL agent (replotted as the semi-transparent line), whereas it remains at zero in DrM.
@@ 4 Experiments
In this section, we present a comprehensive empirical evaluation of MENTOR. Our experimental setup consists of two parts. In Section 4.1, we demonstrate the effectiveness of our method across three simulation benchmarks: DeepMind Control Suite (DMC) (Tassa et al., 2018), Meta-World (MW) (Yu et al., 2020b), and Adroit (Rajeswaran et al., 2017). These benchmarks feature rich visual features and complex dynamics, demanding fine-grained control. Our method consistently outperforms leading visual RL algorithms across these domains. Moreover, one critical limitation in visual RL research is the over-reliance on simulated environments, which raises concerns about the practical applicability of such methods. To mitigate this gap, in Section 4.2, we go beyond simulations and validate the effectiveness of MENTOR in real-world settings on three challenging robotic learning tasks, highlighting the importance of real-world testing.
@@ 4.1 Simulation Experiments
Experimental Settings: We evaluate MENTOR on a diverse set of tasks across three simulation environments with complex dynamics and even sparse reward. The DMC includes challenging tasks like Dog Stand, Dog Walk, Manipulator Bring Ball, and Acrobot Swingup (Sparse), focusing on long-horizon continuous locomotion and manipulation challenges. The MW environment provides a suite of robotic tasks including Assembly, Disassemble, Pick Place, Coffee Push (Sparse), Soccer (Sparse), and Hammer (Sparse), which test the agent’s manipulation abilities and require sequential reasoning. The Adroit environment includes complex robotic manipulation tasks such as Door and Hammer, which involve controlling dexterous hands to interact with articulated objects. Notably, DMC tasks are evaluated using episode reward, while tasks in MW and Adroit are assessed based on success rate. For each method on each task, we conducted experiments using four random seeds; detailed hyperparameters and training settings are provided in Appendix B.
@@
Results: Figure 6 presents performance comparisons between MENTOR and the baselines. In the DMC tasks, Dog Stand and Dog Walk feature high action dimensionality with a 38-dimensional action space representing joint controls for the dog model. These tasks also have complex kinematics involving intricate joint coordination, muscle dynamics, and collision handling, making them challenging to optimize. Our method outperforms the top baseline, achieving approximately 17% and 10% higher episode rewards, respectively. In the MW tasks, the Hammer (Sparse) task stands out. It requires a robotic arm to hammer a nail into a wall, with highly sparse rewards: success yields significantly larger rewards than merely touching or missing the nail. In fact, the reward for failure is only one-thousandth of the success reward, making the task extremely sparse. However, our task-oriented perturbation effectively captures these sparse rewards, reducing the required training frames by 70% compared to the best baseline. In the Adroit tasks, our method achieves nearly 100% success with significantly less training time, while the most competitive counterpart (DrM) requires more frames, and other baselines fail to match performance even after 6 million frames. A key highlight is the Door task, which involves multiple stages of dexterous hand manipulation—grasping, turning, and opening the door. Leveraging the MoE architecture, our method reduces training time to achieve over 80% success by approximately 23% compared to the best baseline. In summary, MENTOR demonstrates superior efficiency and performance compared to the strongest existing model-free visual RL baselines across all 12 tasks,.
@@
Figure 6: Performance comparisons in simulations. This figure compares the performance of our method to DrM, DrQ-v2, ALIX, and TACO across 12 tasks with four random seeds in three different benchmarks. The shaded region indicates standard deviation in DMC and the range of success rates in MW and Adroit.
@@ 4.2 Real-World Experiments
Our real-world RL experiments evaluate the practical applicability of MENTOR in robotic manipulation tasks. We design three tasks to highlight key challenges in real-world robotics: multi-task learning, multi-stage deformable object manipulation, and dynamic skill acquisition.
Experimental Settings: All tasks use a Franka Panda arm for execution and RealSense D435 cameras for RGB visual observations, which include both overall and close-up views to capture global and local information. The reward functions are based on the absolute distance between the current and desired states. To prevent trajectory overfitting, the end-effector’s initial position is randomly sampled from a predefined region at the start of each episode. Tasks are described below and shown in Figure 7. Further details can be found in Appendix C.
Peg Insertion: This task simulates an assembly-line scenario where fine-grained insertion of various objects are required. The agent needs to develop multi-task learning skills to insert pegs with three different shapes (Star, Triangle, and Arrow) into corresponding sockets. Training such agents in simulators is difficult due to the complexities of contact-rich interactions, making this task ideal for real-world reinforcement learning and evaluation.
Cable Routing: Manipulating deformable cables presents significant challenges due to the complexities of modeling and simulating their physical dynamics, making this task ideal for direct, model-free visual RL training in real-world environments. In this scenario, the robot must guide a cable into two parallel slots. Since both slots cannot be filled simultaneously, the agent must perform the task sequentially, requiring long-horizon, multi-stage planning to successfully accomplish the task.
@@
Figure 7: Real-world experiments (up to down rows: Peg Insertion, Cable Routing, and Table Golf). This set of images illustrates the execution of the learned visual policy trained using MENTOR. The agent consistently and accurately accomplishes tasks even in the presence of human disturbances.
Results: Our policies demonstrate robust performance during evaluation as shown in Figure 7. In Peg Insertion, the agent randomly selects a peg from the shelf and inserts it from varying initial positions. It gradually learns to align the peg shape with the corresponding hole and adjust the angle for accurate insertion. During one execution, as the peg nears the hole, we manually disturb by altering the robot arm’s pose significantly. Despite this interference, the agent successfully completes the task relying solely on visual observations. In Cable Routing, where the cable cannot be placed into parallel slots simultaneously, the agent learns to prioritize routing it into the farther slot first, then into the closer one. This second step requires careful handling to avoid dislodging the cable from the first slot. During execution, if the cable is randomly removed from the slot, the agent can visually detect this issue and re-route it back into position. In Tabletop Golf, the agent must master two key skills: striking the ball with the correct direction and force, and repositioning the club to follow the ball after the strike. Due to a "water hazard", the ball cannot be struck directly toward the target hole from its starting position. The agent learns to angle its shots to bypass the hazard and guide the ball into the hole. No interference is applied during this task, as the ball’s rolling on the grass-like surface introduces sufficient variability.
@@
Furthermore, the results confirm the effectiveness of our technical contributions. When the MoE structure is removed from the agent (i.e., replaced with an MLP, as in MENTOR w/o MoE), overall performance drops by nearly 30%. Additionally, further switching the task-oriented perturbation mechanism to basic random perturbation (as in DrM) leads to an additional performance decline of approximately 30%. We further extend the training process of the DrM baseline to reach the same performance level as MENTOR, with the training time comparison shown in Figure 8, which demonstrates an average 37% improvement in time efficiency for our method. These findings underscore the importance of each component in achieving superior results.
@@ 5 Related work
Visual reinforcement learning (RL), which operates on pixel observations rather than ground-truth state vectors, faces significant challenges in decision-making due to the high-dimensional nature of visual inputs and the difficulty in extracting meaningful features for policy optimization (Ma et al., 2022; Choi et al., 2023; Ma et al., 2022). Despite these challenges, there has been considerable progress in this area. Methods such as Hafner et al. (2019; 2020; 2023); Hansen et al. (2022) improve visual RL by building world models. Other approaches (Yarats et al., 2021; Kostrikov et al., 2020; Laskin et al., 2020b), use data augmentation to enhance learning robustness from pixel inputs. Contrastive learning, as in Laskin et al. (2020a); Zheng et al. (2023), aids in learning more informative state and action representations. Additionally, Cetin et al. (2022) applies regularization to prevent catastrophic self-overfitting, while DrM (Xu et al., 2023) enhances exploration by periodically perturbing the agent’s parameters. Despite recent progress, these methods still suffer from low sample efficiency in complex robotic tasks. In this paper, we propose enhancing the agent’s learning capability by replacing the standard MLP backbone with an MoE architecture. This dynamic expert learning mechanism helps mitigate gradient conflicts in complex scenarios.
@@ 6 Conclusion
In this paper, we present MENTOR, a state-of-the-art model-free visual RL framework that achieves superior performance in challenging robotic control tasks. MENTOR enhances learning efficiency through two key improvements in both agent network architecture and optimization. Replacing the traditional multi-layer perceptron (MLP) backbone with mixture-of-experts (MoE) structure enables the agent to dynamically allocate learning gradients to modular experts via a sparse routing system, mitigating gradient conflicts in complex scenarios. Additionally, the introduction of a task-oriented perturbation mechanism helps refine the agent’s weights toward optimal solutions by sampling perturbation candidates from a heuristically updated distribution formed by the top-performing agents during RL training. MENTOR consistently outperforms the strongest baselines across 12 tasks in three simulation benchmark environments. Furthermore, we extend our evaluation beyond simulations, demonstrating the effectiveness of MENTOR in real-world settings on three challenging self-designed robotic manipulation tasks, which highlights its sample efficiency and practical value. We believe MENTOR is a capable visual RL algorithm with the potential to push the boundaries of RL application in real-world robotic tasks.
@@ Appendix A Algorithm Details
1:Input: Batch of visual inputs $\{\mathbf{x}_{b}\}_{b=1}^{B}$
2:Output: Final actions $\{a_{b}\}_{b=1}^{B}$, Load balancing loss $\mathcal{L}_{\text{LB}}$
3:Initialize experts $\{\operatorname{FFN}_{1},\operatorname{FFN}_{2},\dots,\operatorname{FFN}_{N}\}$
4:$\{\mathbf{z}_{b}\}_{b=1}^{B}\leftarrow\operatorname{Encoder}(\{\mathbf{x}_{b}\}_{b=1}^{B})$
5:$\{\mathbf{h}_{b}\}_{b=1}^{B}\leftarrow h(\{\mathbf{z}_{b}\}_{b=1}^{B})$
6:for each input $b=1$ to $B$ do
7:     $\mathcal{E}_{b}\leftarrow\operatorname{topk}(\mathbf{h}_{b},k)$
8:     $w_{b}(i)\leftarrow\operatorname{softmax}(\mathbf{h}_{b,\mathcal{E}_{b}})[i],\quad\forall i\in\mathcal{E}_{b}$
9:     for each expert $i\in\mathcal{E}_{b}$ do
10:         $\mathbf{f}_{b,i}\leftarrow\operatorname{FFN}_{i}(\mathbf{z}_{b})$
11:     end for
12:     $a_{b}\leftarrow\operatorname{ActionProjector}\left(\sum_{i\in\mathcal{E}_{b}}w_{b}(i)\,\mathbf{f}_{b,i}\right)$
13:end for
14:// Compute load balancing loss
15:$\{\mathbf{pr}_{b}\}_{b=1}^{B}\leftarrow\operatorname{softmax}(\{\mathbf{h}_{b}\}_{b=1}^{B})$
16:$p(i)\leftarrow\frac{1}{B}\sum_{b=1}^{B}pr_{b}(i),\quad\forall i$
17:$\mathcal{L}_{\text{LB}}\leftarrow-H(p)=\sum_{i=1}^{N}p(i)\log(p(i))$

### v3

@@ 3.2 Task-oriented Perturbation Mechanism
Figure 5: Validation of task-oriented perturbation on Hopper Hop (a MENTOR, DrM, and DrQ-v2 agent trained on the Hopper Hop task during the first 1M frames). Our method consistently achieves higher episode rewards with a consistently lower dormant ratio. (c) shows the episode reward obtained by perturbation candidate $\phi$ sampled from $\Phi_{\text{oriented}}$ steadily increases and occasionally surpasses that of the corresponding RL agent (replotted as the light red line), whereas in DrM, the reward remains at zero due to the use of randomly generated perturbation parameters.
@@ 4 Experiments
In this section, we present a comprehensive empirical evaluation of MENTOR. Section 4.1 showcases its effectiveness on three simulation benchmarks: DeepMind Control Suite (DMC) (Tassa et al., 2018), Meta-World (MW) (Yu et al., 2020b), and Adroit (Rajeswaran et al., 2017), which feature rich visuals and complex dynamics. MENTOR consistently outperforms leading visual RL algorithms. However, a critical limitation in visual RL research is the over-reliance on simulated environments, raising concerns about practical applicability. To bridge this gap, Section 4.2 validates MENTOR on three challenging real-world robotic tasks, highlighting the importance of real-world testing.
@@ 4.1 Simulation Experiments
Experimental Settings: We evaluate MENTOR on a diverse set of tasks across three simulation environments with complex dynamics and even sparse reward. The DMC includes challenging tasks like Dog Stand, Dog Walk, Manipulator Bring Ball, and Acrobot Swingup (Sparse), focusing on long-horizon continuous locomotion and manipulation challenges. The MW environment provides a suite of robotic tasks including Assembly, Disassemble, Pick Place, Coffee Push (Sparse), Soccer (Sparse), and Hammer (Sparse), which test the agent’s manipulation abilities and require sequential reasoning. The Adroit environment includes complex robotic manipulation tasks such as Door and Hammer, which involve controlling dexterous hands to interact with articulated objects. Notably, DMC tasks are evaluated using episode reward, while tasks in MW and Adroit are assessed based on success rate. We conducted experiments with four random seeds on each task, with detailed hyperparameters and training settings provided in Appendix B.
@@
Results: Figure 6 presents performance comparisons between MENTOR and the baselines. In the DMC tasks, Dog Stand and Dog Walk feature high action dimensionality with a 38-dimensional action space representing joint controls for the dog model. These tasks also have complex kinematics involving intricate joint coordination, muscle dynamics, and collision handling, making them challenging to optimize. Our method outperforms the top baseline, achieving approximately 17% and 10% higher episode rewards, respectively. In the MW tasks, the Hammer (Sparse) task stands out. It requires a robotic arm to hammer a nail into a wall, with highly sparse rewards: success yields significantly larger rewards than merely touching or missing the nail. In fact, the reward for failure is only one-thousandth of the success reward, making the task extremely sparse. However, our task-oriented perturbation effectively captures these sparse rewards, reducing the required training frames by 70% compared to the best baseline. In the Adroit tasks, our method achieves nearly 100% success with significantly less training time, while the most competitive counterpart (DrM) requires more frames, and other baselines fail to match performance even after 6 million frames. A key highlight is the Door task, which involves multiple stages of dexterous hand manipulation—grasping, turning, and opening the door. Leveraging the MoE architecture, our method reduces training time to achieve over 80% success by approximately 23% compared to the best baseline. In summary, MENTOR demonstrates superior efficiency and performance compared to the strongest existing model-free visual RL baselines across all 12 tasks. For the robustness against disturbances of agent trained by MENTOR, please see Appendix I.
Ablation Study: We conduct a detailed ablation study to demonstrate the significance contribution of MoE and Task-oriented Perturbation separately in Appendix C.
@@
Figure 6: Performance comparisons in simulations. This figure compares the performance of our method to DrM, DrQ-v2, ALIX, and TACO across 12 tasks with four random seeds in three different benchmarks (DMC, MW, and Adroit). The shaded region indicates standard deviation in DMC and the range of success rates in MW and Adroit.
@@ 4.2 Real-World Experiments
Our real-world RL experiments evaluate MENTOR on three key challenges: multi-task learning, multi-stage deformable object manipulation, and dynamic skill acquisition with policies trained from scratch in the real world.
Experimental Settings: We use a Franka Panda arm for execution and RealSense D435 cameras for RGB observations, capturing both global and local views. Rewards are based on the absolute distance between current and desired states. To prevent trajectory overfitting, the end-effector’s initial position is randomly sampled within a predefined region at the start of each episode. Tasks are detailed below, with illustrations in Figure 7. Additional details are in Appendix D.
Peg Insertion: This task mimics assembly-line scenarios requiring the agent to insert pegs of three shapes (Star, Triangle, Arrow) into corresponding sockets. Simulating contact-rich interactions in these multi-task settings is highly challenging, making real-world evaluation essential.
Cable Routing: Manipulating deformable cables presents significant challenges due to the complexities of modeling and simulating their physical dynamics, making this task ideal for direct, model-free visual RL training in real-world environments. The robot must guide a deformable cable sequentially into two parallel slots. Since both slots cannot be filled simultaneously, the agent must perform the task sequentially, requiring long-horizon, multi-stage planning to successfully accomplish the task.
@@
Figure 7: Real-world experiments (up to down rows: Peg Insertion, Cable Routing, and Tabletop Golf). This set of images illustrates the execution of the learned visual policy trained using MENTOR. The agent consistently and accurately accomplishes tasks even in the presence of human disturbances.
Results: Our policies exhibit strong performance during evaluation, as shown in Figure 7. In Peg Insertion, the agent randomly picks a peg and inserts it from varying initial positions, learning to align the peg with the hole and adjust the angle for accurate insertion. During one execution, as the peg nears the hole, we manually disturb by altering the robot arm’s pose significantly. Despite this interference, the agent successfully completes the task relying solely on visual observations. In Cable Routing, the agent learns to prioritize routing it into the farther slot first, then into the closer one. This second step requires careful handling to avoid dislodging the cable from the first slot. During execution, if the cable is randomly removed from the slot, the agent can visually detect this issue and re-route it back into position. In Tabletop Golf, the agent must master two key skills: striking the ball with the correct direction and force, and repositioning the club to follow the ball after the strike. Due to a “water hazard”, the ball cannot be struck directly toward the target hole from its starting position. The agent learns to angle its shots to bypass hazards and guide the ball into the hole, even as the ball’s rolling on the grass-like surface naturally introduces significant variability.
@@
Furthermore, the results confirm the effectiveness of our technical contributions. When the MoE structure is removed from the agent (i.e., replaced with an MLP, as in MENTOR w/o MoE), overall performance drops by nearly 30%. Additionally, further switching the task-oriented perturbation mechanism to basic random perturbation (as in DrM) leads to an additional performance decline of approximately 30%.
@@ 5 Related work
Visual reinforcement learning (RL), which operates on pixel observations rather than ground-truth state vectors, faces significant challenges in decision-making due to the high-dimensional nature of visual inputs and the difficulty in extracting meaningful features for policy optimization (Ma et al., 2022; Choi et al., 2023; Ma et al., 2022). Despite these challenges, there has been considerable progress in this area. Methods such as Hafner et al. (2019, 2020, 2023); Hansen et al. (2022) improve visual RL by building world models. Other approaches (Yarats et al., 2021; Kostrikov et al., 2020; Laskin et al., 2020a), use data augmentation to enhance learning robustness from pixel inputs. Contrastive learning, as in Laskin et al. (2020b); Zheng et al. (2023), aids in learning more informative state and action representations. Additionally, Cetin et al. (2022) applies regularization to prevent catastrophic self-overfitting, while DrM (Xu et al., 2023) enhances exploration by periodically perturbing the agent’s parameters. Despite recent progress, these methods still suffer from low sample efficiency in complex robotic tasks. In this paper, we propose enhancing the agent’s learning capability by replacing the standard MLP backbone with an MoE architecture. This dynamic expert learning mechanism helps mitigate gradient conflicts in complex scenarios.
@@ 6 Conclusion
In this paper, we present MENTOR, a state-of-the-art model-free visual RL framework that achieves superior performance in challenging robotic control tasks. MENTOR enhances learning efficiency through two key improvements in both agent network architecture and optimization. MENTOR consistently outperforms the strongest baselines across 12 tasks in three simulation benchmark environments. Furthermore, we extend our evaluation beyond simulations, demonstrating the effectiveness of MENTOR in real-world settings on three challenging robotic manipulation tasks. We believe MENTOR is a capable visual RL algorithm with the potential to push the boundaries of RL application in real-world robotic tasks. While MENTOR has proven effective, it has been evaluated on single tasks with a single robot embodiment. Future work could scale our method to handle hundreds of tasks or diverse robot embodiments, enabling broader real-world applications.
Impact Statement
This paper presents work whose goal is to advance the field of Model-free Visual Reinforcement Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Algorithm Details
0:  Batch of visual inputs $\{\mathbf{x}_{b}\}_{b=1}^{B}$
0:  Final actions $\{a_{b}\}_{b=1}^{B}$, Load balancing loss $\mathcal{L}_{\text{LB}}$
  Initialize experts $\{\operatorname{FFN}_{1},\operatorname{FFN}_{2},\dots,\operatorname{FFN}_{N}\}$
  $\{\mathbf{z}_{b}\}_{b=1}^{B}\leftarrow\operatorname{Encoder}(\{\mathbf{x}_{b}\}_{b=1}^{B})$
  $\{\mathbf{h}_{b}\}_{b=1}^{B}\leftarrow h(\{\mathbf{z}_{b}\}_{b=1}^{B})$
  for $b=1$ to $B$ do
     $\mathcal{E}_{b}\leftarrow\operatorname{topk}(\mathbf{h}_{b},k)$
     $w_{b}(i)\leftarrow\operatorname{softmax}(\mathbf{h}_{b,\mathcal{E}_{b}})[i]$, $\forall i\in\mathcal{E}_{b}$
     for each $i\in\mathcal{E}_{b}$ do
        $\mathbf{f}_{b,i}\leftarrow\operatorname{FFN}_{i}(\mathbf{z}_{b})$
     end for
     $a_{b}\leftarrow\operatorname{ActionProjector}\left(\sum_{i\in\mathcal{E}_{b}}w_{b}(i)\,\mathbf{f}_{b,i}\right)$
  end for
  // Compute load balancing loss
  $\{\mathbf{pr}_{b}\}_{b=1}^{B}\leftarrow\operatorname{softmax}(\{\mathbf{h}_{b}\}_{b=1}^{B})$
  $p(i)\leftarrow\frac{1}{B}\sum_{b=1}^{B}pr_{b}(i)$, $\forall i$
  $\mathcal{L}_{\text{LB}}\leftarrow-H(p)=\sum_{i=1}^{N}p(i)\log(p(i))$
Figure 8: Ablation study on the hyperparameter of MoE. We evaluated MENTORon MW Hammer (Sparse) wit different MoE settings.
@@ Appendix B Simulation Experimental Settings
For the hyperparameters used in MoE module, Figure 8 shows the ablation study on the number of experts and top_k chosen experts. The results indicate the optimal setting for the Hammer task is MoE has around 8 experts, performance remains consistent across 4, 8, and 32 experts as long as top_k = 4. That is, given a proper top_k, the performance is not sensitive to the number of experts.
Figure 9: Ablation study on key contributions. This figure shows the experiment result of four ablated versions of MENTOR using the same four random seeds as in the original experiments, illustrating our Mixture-of-Experts (MoE) and Task-oriented Perturbation (TP) are both significant to improving performance.
Table 2: Sample efficiency comparison across tasks.
Sample Efficiency (Training Time) Hopper Hop Disassemble Coffee Push Hammer
MENTOR (Ours) 0.6167 0.7056 0.8066 0.7167
MENTOR_w/o_TP 1 0.8505 0.9481 0.875
MENTOR_w/o_MoE 0.85 1 1 1
Appendix C Ablation Study on Key Contributions
We conducted additional ablation studies on four diverse tasks: Hopper Hop, Disassemble, Coffee-Push (Sparse), and Hammer (Sparse). These studies aim to decouple the effects of the MoE architecture and the Task-oriented Perturbation (TP) mechanism proposed in our paper.
For the experiments, we evaluate four ablated versions of MENTOR using the same four random seeds as in the original experiments, as shown in Figure 9:
  • MENTOR: Full model with both MoE and Task-oriented Perturbation.
  • MENTOR_w/o_TP: Task-oriented Perturbation is replaced with random perturbation.
  • MENTOR_w/o_MoE: The policy backbone uses an MLP architecture instead of MoE.
  • MENTOR_w/o_TP_MoE: Neither MoE nor Task-oriented Perturbation is used.
The results, summarized below, demonstrate the individual contributions of each component:
The overall sample efficiency and performance of MENTOR_w/o_TP and MENTOR_w/o_MoE remain lower than the full MENTOR model. This underscores the complementary nature of these two components in enhancing the overall learning efficiency and robustness of MENTOR.
Considering the performance, MENTOR_w/o_MoE and MENTOR_w/o_TP also consistently outperform MENTOR_w/o_TP_MoE, indicating that both the MoE architecture and Task-oriented Perturbation independently contribute to improved policy learning.
Considering efficiency, the following part shows a quantitative result based on standard training time.
Standard Training Time: Let $T_{\text{MENTOR}}$, $T_{\text{MENTOR\_w/o\_MOE}}$, and $T_{\text{MENTOR\_w/o\_TP}}$ denote the time required for the three different methods to reach the same performance (the final performance of the worst method). The standard training time $T_{\text{standard}}$ is defined as the training time for the worst method to achieve this performance:
$T_{\text{standard}}=\max(T_{\text{MENTOR}},T_{\text{MENTOR\_w/o\_MOE}},T_{\text{MENTOR\_w/o\_TP}})$
We define normalized sample efficiency as $\frac{T_{*}}{T_{\text{standard}}}$ (lower is better).
Table 2 shows the normalized sample efficiency of each setting. MENTOR (Ours) achieves an average of 22.6% and 26.1% less training time over the 4 tasks compared with MENTOR_w/o_TP and MENTOR_w/o_MoE.